# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 3. Local well-posedness
Now, we invoke the $\mathcal{Z}^{2}_{\rho}$ bound (3.5). First, choose $n_{1}$ as $2^{n_{1}}=C(1+M_{0}^{2})$ with $C\gg_{k}1$, so that
$|\mathcal{I}_{k}(t)|\leq\frac{1}{2}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},\qquad t\in[0,T_{0}].$
Next, we have the modified energy inequality
$\frac{d}{dt}\big(\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)\big)\lesssim_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}_{\rho}(\phi)+\big[\mathcal{E}^{2}_{\rho}(\phi)\big]^{2}\Big)\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}.$
Noticing that $\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}\lesssim\mathcal{E}^{1}_{\rho}(u^{(\ell)}(t))+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\lesssim\mathcal{E}^{2}_{\rho}(\phi)+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}$, we see that the function
$X_{k}(t):=\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)+\mathcal{E}^{2}_{\rho}(\phi)$
@@
$X_{k}^{\prime}(t)\leq C_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}_{\rho}(\phi)+\big[\mathcal{E}^{2}_{\rho}(\phi)\big]^{2}\Big)X_{k}(t),\qquad t\in[0,T_{0}].$
Integrating it and recalling the definition of $T_{0}$, we obtain
$\|X_{k}\|_{L^{\infty}_{T_{0}}}\leq X_{k}(0)e^{C_{k}T_{0}(1+\delta^{-1})\big(\mathcal{E}^{2}_{\rho}(\phi)+[\mathcal{E}^{2}_{\rho}(\phi)]^{2}\big)}\leq Ce^{C_{k}}\Big(\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}_{\rho}(\phi)\Big).$
Since $\mathcal{E}^{k}_{\rho}(u^{(\ell)}(t))\lesssim X_{k}(t)$ for $t\in[0,T_{0}]$ by (3.5), this yields claim (3.6). ∎
Proposition 3.3 ($\mathcal{Z}^{1}_{\rho}$ difference estimate).
Let $\delta>0$, $\rho>0$, $M_{0}>0$, and let $\phi,\psi\in\mathcal{Z}_{\rho}^{2}$ satisfy $\mathcal{E}_{\rho}^{2}(\phi)\leq M_{0}^{2}$ and $\mathcal{E}_{\rho}^{2}(\psi)\leq M_{0}^{2}$. For $\ell,m\geq 0$, let $u^{(\ell)},v^{(m)}\in C([0,T_{0}];\mathcal{Z}^{\infty}_{\rho})$ be the solutions to (3.1) with the mollification parameters $\ell,m$ and the initial data $P_{\leq\ell}\phi$ and $P_{\leq m}\psi$, respectively, where $T_{0}=T_{0}(\delta,M_{0})\in(0,1]$ is the constant given in Proposition 3.2. Then, there exists $T_{*}\in(0,T_{0}]$ with $T_{*}\gtrsim\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ such that we have
(3.15) $\|d_{\rho}^{1}(u^{(\ell)},v^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0})d_{\rho}^{1}(\phi,\psi)+(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
@@
(3.16) $\|d_{\rho}^{1}(u^{(\ell)},u^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
@@
Take an arbitrary $T\in(0,T_{0}]$, and recall $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$ by the definition of $T_{0}$. We see from Proposition 3.2 and (2.9) that
(3.17) $\|\mathcal{E}^{2}_{\rho}(\zeta_{\ell})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}_{\rho}(\eta_{m})\|_{L^{\infty}_{T}}\lesssim\|\mathcal{E}^{2}_{\rho}(u^{(\ell)})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}_{\rho}(v^{(m)})\|_{L^{\infty}_{T}}\lesssim M_{0}^{2}.$
@@
$\|d^{1}_{\rho}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}$ $\lesssim(1+M_{0})d^{1}_{\rho}(P_{\leq\ell}\phi,P_{\leq m}\psi)$
$\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3}),$
@@
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{3})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}$
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{4})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}.$
Combining the above estimates and (2.8), we obtain
$\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}L^{2}})\|d^{1}_{\rho}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}])\|_{L^{\infty}_{T}L^{2}}$
@@
$\lesssim\|d^{1}_{\rho}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}+(1+M_{0})\|\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{2}}$
$\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}$
$\quad+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{5})+T(1+\delta^{-1})(1+M_{0}^{3})\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}},$
where at the second inequality we have used the equation as $\Phi^{(\ell)}[u^{(\ell)}]=u^{(\ell)}-U(t)P_{\leq\ell}\phi$. Imposing further smallness $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$ if necessary, we can absorb the last term into the left-hand side and obtain the control of lower order terms,
(3.18) $\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}.$
We move on to the energy estimate for the highest order term. Using the equation (3.1),
$\frac{d}{dt}\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}$ $=2\operatorname{Re}\big({-}i\partial_{x}\big(P_{\leq\ell}\mathcal{N}(\zeta_{\ell}(t))-P_{\leq m}\mathcal{N}(\eta_{m}(t))\big),\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big)_{L^{2}}$
We first rearrange the frequency projections by using Bernstein’s inequality, integration by parts, Lemma 2.1 and the uniform $\mathcal{Z}^{2}_{\rho}$ bound (3.17), as
@@
Next, we expand the difference of the nonlinearity as follows:
@@
The contribution of the $\mathcal{N}_{\delta}$ part is easily estimated with Lemma 2.1, by
$\|\partial_{x}(\mathcal{N}_{\delta}(\zeta_{\ell})-\mathcal{N}_{\delta}(\eta_{m}))\|_{L^{\infty}_{T}L^{2}}\|\partial_{x}\omega_{\ell,m}\|_{L^{\infty}_{T}L^{2}}\lesssim\delta^{-1}(1+M_{0}^{2})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
For the other terms, integration by parts argument with Lemma 2.2 shows that the only problematic interaction is
$2\beta\operatorname{Re}\big(\mu_{\ell,m}(t)(iH)(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{\omega_{\ell,m}(t)})),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}},$
while the others can be bounded by $M_{0}^{2}\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}$. So far, we have proven that
@@
$\quad+C(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{4}+C(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
(3.21) $|\mathcal{R}_{2}(t)|$ $\lesssim(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{6}+(1+\delta^{-1})(M_{0}^{2}+M_{0}^{4})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
$|\mathcal{R}_{3}(t)|$ $\lesssim(1+2^{n_{2}})M_{0}^{2}\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\quad+(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\|\mathcal{J}(t)\|_{L^{\infty}_{T}}\lesssim 2^{-n_{2}}M_{0}^{2}\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
@@
$\big\|\partial_{x}\big(u^{(\ell)}-v^{(m)}\big)\big\|_{L^{\infty}_{T}L^{2}}^{2}$ $\lesssim\big\|\partial_{x}\big(P_{\leq\ell}\phi-P_{\leq m}\psi\big)\big\|_{L^{2}}^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}+2^{-n_{2}}M_{0}^{2}\|d^{1}_{\rho}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\lesssim\Big(d^{1}_{\rho}(\phi,\psi)+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}\Big(\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}.$
Combining it with (3.18), we come to
$\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\big[(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})\big]^{\frac{1}{2}}$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}^{\frac{1}{2}}\Big(\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big).$
Now, we choose $n_{2}$ as $2^{n_{2}}=C(1+M_{0}^{2})$ for a large constant $C>0$, and impose further smallness $T\ll\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ if necessary, so that the coefficient of $\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$ on the right-hand side becomes sufficiently small. Then, this term is absorbed into the left-hand side and we finally obtain
$\|d^{1}_{\rho}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi)+(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m})(1+M_{0}^{3}),$
@@
At this stage, we are now ready to complete the proof of local well-posedness of Theorem 1.1.

### v2

@@ 3. Local well-posedness
Now, we invoke the $\mathcal{Z}^{2}$ bound (3.5) and choose $n_{1}$ as $2^{n_{1}}=C(1+M_{0}^{2})$ with $C\gg_{k}1$. Then, we have
$|\mathcal{I}_{k}(t)|\leq\frac{1}{2}\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2},\qquad t\in[0,T_{0}],$
and the modified energy inequality
$\frac{d}{dt}\big(\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)\big)\lesssim_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}(\phi)+\big[\mathcal{E}^{2}(\phi)\big]^{2}\Big)\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2},\qquad t\in[0,T_{0}].$
Noticing that $\|\partial_{x}u^{(\ell)}(t)\|_{H^{k-1}}^{2}\lesssim\mathcal{E}^{1}(u^{(\ell)}(t))+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}\lesssim\mathcal{E}^{2}(\phi)+\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}$, we see that the function
$X_{k}(t):=\|\partial_{x}^{k}u^{(\ell)}(t)\|_{L^{2}}^{2}+\mathcal{I}_{k}(t)+\mathcal{E}^{2}(\phi)$
@@
$\frac{d}{dt}X_{k}(t)\leq C_{k}(1+\delta^{-1})\Big(\mathcal{E}^{2}(\phi)+\big[\mathcal{E}^{2}(\phi)\big]^{2}\Big)X_{k}(t),\qquad t\in[0,T_{0}].$
Using Gronwall’s inequality and recalling $T_{0}(1+\delta^{-1})((1+2^{n_{0}})M_{0}^{2}+M_{0}^{4})\ll 1$, we obtain
$\|X_{k}\|_{L^{\infty}_{T_{0}}}\leq X_{k}(0)e^{C_{k}T_{0}(1+\delta^{-1})\big(\mathcal{E}^{2}(\phi)+[\mathcal{E}^{2}(\phi)]^{2}\big)}\leq Ce^{C_{k}}\Big(\|\partial_{x}^{k}P_{\leq\ell}\phi\|_{L^{2}}^{2}+\mathcal{E}^{2}(\phi)\Big).$
Since $\mathcal{E}^{k}(u^{(\ell)}(t))\lesssim X_{k}(t)$ for $t\in[0,T_{0}]$ by (3.5), this yields (3.6). ∎
Proposition 3.3 ($\mathcal{Z}^{1}$ difference estimate).
Let $\delta>0$, $M_{0}>0$, and let $\phi,\psi\in\mathcal{Z}^{2}$ satisfy $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$ and $\mathcal{E}^{2}(\psi)\leq M_{0}^{2}$. For $\ell,m\geq 0$, let $u^{(\ell)},v^{(m)}\in C([0,T_{0}];\mathcal{Z}^{\infty})$ be the solutions to (3.1) with the mollification parameters $\ell,m$ and the initial data $P_{\leq\ell}\phi$ and $P_{\leq m}\psi$, respectively, where $T_{0}\in(0,1]$ is the constant given in Proposition 3.2. Then, there exists $T_{*}\in(0,T_{0}]$ with $T_{*}\gtrsim\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ such that we have
(3.15) $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
@@
(3.16) $\|d^{1}(u^{(\ell)},u^{(m)})\|_{L_{T_{*}}^{\infty}}\lesssim(1+M_{0}^{3})(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m}).$
@@
Take an arbitrary $T\in(0,T_{0}]$ with $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$. We see from Proposition 3.2 and (2.9) that
(3.17) $\|\mathcal{E}^{2}(\zeta_{\ell})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}(\eta_{m})\|_{L^{\infty}_{T}}\lesssim\|\mathcal{E}^{2}(u^{(\ell)})\|_{L^{\infty}_{T}}+\|\mathcal{E}^{2}(v^{(m)})\|_{L^{\infty}_{T}}\lesssim M_{0}^{2}.$
@@
$\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}$ $\lesssim(1+M_{0})d^{1}(P_{\leq\ell}\phi,P_{\leq m}\psi)$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3}),$
@@
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{3})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}$
$\lesssim(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{4})+T(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}.$
Combining (2.8) and the above estimates, we obtain
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+\|\Phi^{(\ell)}[u^{(\ell)}]\|_{L^{\infty}_{T}L^{2}})\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}])\|_{L^{\infty}_{T}L^{2}}$
@@
$\lesssim\|d^{1}(U(t)P_{\leq\ell}\phi,U(t)P_{\leq m}\psi)\|_{L^{\infty}_{T}}+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}+(1+M_{0})\|\Phi^{(\ell)}[u^{(\ell)}]-\Phi^{(m)}[v^{(m)}]\|_{L^{\infty}_{T}L^{2}}$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}$
$\quad+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(1+M_{0}^{5})+T(1+\delta^{-1})(1+M_{0}^{3})\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}},$
where at the second inequality we have used the equation as $\Phi^{(\ell)}[u^{(\ell)}]=u^{(\ell)}-U(t)P_{\leq\ell}\phi$. Imposing further smallness $T(1+\delta^{-1})(1+M_{0}^{4})\ll 1$ if necessary, we can absorb the last term into the left-hand side and obtain
(3.18) $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\|\partial_{x}(u^{(\ell)}-v^{(m)})\|_{L^{\infty}_{T}L^{2}}.$
It remains to estimate the last term on the right-hand side of (3.18). Using the equation (3.1), we see that
$\frac{d}{dt}\big\|\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big\|_{L^{2}}^{2}$ $=2\operatorname{Re}\big({-}i\partial_{x}\big(P_{\leq\ell}\mathcal{N}(\zeta_{\ell}(t))-P_{\leq m}\mathcal{N}(\eta_{m}(t))\big),\partial_{x}\big(u^{(\ell)}(t)-v^{(m)}(t)\big)\big)_{L^{2}}.$
We first rearrange the frequency projections by using Bernstein’s inequality, integration by parts, Lemma 2.1 and (3.17) as
@@
Next, we expand the difference of the nonlinearity as
@@
The contribution of the $\mathcal{N}_{\delta}$ part is easily estimated with Lemma 2.1 by
$\|\partial_{x}(\mathcal{N}_{\delta}(\zeta_{\ell})-\mathcal{N}_{\delta}(\eta_{m}))\|_{L^{\infty}_{T}L^{2}}\|\partial_{x}\omega_{\ell,m}\|_{L^{\infty}_{T}L^{2}}\lesssim\delta^{-1}(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
For the remaining terms, since $\omega_{\ell,m}$ corresponds to the difference $u^{(\ell)}-v^{(m)}$, by integration by parts, Lemma 2.2 and (3.17), we see that the only problematic interaction is
$2\beta\operatorname{Re}\big(\mu_{\ell,m}(t)(iH)(\mu_{\ell,m}(t)\partial_{x}^{2}\overline{\omega_{\ell,m}(t)})),\partial_{x}\omega_{\ell,m}(t)\big)_{L^{2}}.$
Indeed, the others can be bounded by $M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}$. So far, we have proven that
@@
$\quad+C(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{4}+C(1+\delta^{-1})(1+M_{0}^{2})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
(3.21) $|\mathcal{R}_{2}(t)|$ $\lesssim(2^{-\ell}+2^{-m})(1+\delta^{-1})M_{0}^{6}+(1+\delta^{-1})(M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
$|\mathcal{R}_{3}(t)|$ $\lesssim(1+2^{n_{2}})M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\quad+(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\|\mathcal{J}(t)\|_{L^{\infty}_{T}}\lesssim 2^{-n_{2}}M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2},$
@@
(3.23) $\big\|\partial_{x}\big(u^{(\ell)}-v^{(m)}\big)\big\|_{L^{\infty}_{T}L^{2}}^{2}$ $\lesssim\big\|\partial_{x}\big(P_{\leq\ell}\phi-P_{\leq m}\psi\big)\big\|_{L^{2}}^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}+2^{-n_{2}}M_{0}^{2}\|d^{1}(\zeta_{\ell},\eta_{m})\|_{L^{\infty}_{T}}^{2}.$
@@
$\lesssim\Big(d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}+(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}\Big(\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big)^{2}.$
Combining this with (3.18), we come to
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$
$\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\ell}+2^{-m})(1+M_{0}^{3})+\big[(2^{-\ell}+2^{-m})T(1+\delta^{-1})(M_{0}^{4}+M_{0}^{6})\big]^{\frac{1}{2}}$
$\quad+\Big\{T(1+\delta^{-1})(1+2^{n_{2}}M_{0}^{2}+M_{0}^{4})+2^{-n_{2}}M_{0}^{2}\Big\}^{\frac{1}{2}}\Big(\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}+(2^{-\ell}+2^{-m})(M_{0}+M_{0}^{2})\Big).$
Now, we choose $n_{2}$ as $2^{n_{2}}=C(1+M_{0}^{2})$ with $C\gg 1$, and impose further smallness $T\ll\big[(1+\delta^{-1})(1+M_{0}^{4})\big]^{-1}$ if necessary, so that the coefficient of $\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}$ on the right-hand side becomes sufficiently small. Then, this term is absorbed into the left-hand side and we finally obtain
$\|d^{1}(u^{(\ell)},v^{(m)})\|_{L^{\infty}_{T}}\lesssim(1+M_{0})d^{1}(\phi,\psi)+(2^{-\frac{1}{2}\ell}+2^{-\frac{1}{2}m})(1+M_{0}^{3}),$
@@
At this stage, we are now ready to complete the proof of Theorem 1.1.